# BotArtist: Generic approach for bot detection in Twitter via semi-automatic machine learning pipeline

### v3

@@ Introduction
Having reduced the problem’s dimensionality through feature selection, we now focus on identifying the model and its configuration (hyper-parameter set) that can provide accurate predictions over unseen data samples. We selected three well-known machine learning classification models: SVM (Boser, Guyon, and Vapnik 1992), RandomForest (Breiman 2001), and state-of-the-art XGBoost (Extreme Gradient Boosting) (Chen et al. 2015).
@@
For each model, we defined large hyper-parameter ranges to cover a variety of configurations. To reduce complexity and execution time, we developed a sampling method that randomly selects $C$ unique configurations per model (in our case, $C=50$). This method allows us to estimate different possible configurations without sacrificing execution time. Additionally, since the dataset may be affected by class imbalance, we compute class-specific weights to construct selected classifiers with consideration of class imbalance.
@@
Afterward, we select the best classifier and configuration based on the average validation (F1-score). This model configuration becomes the final model, which will be evaluated over the hold-out (testing) data portion. In the final stage of the model development, we provide model explainability using the SHAP game-theoretical approach (Shapley 1953). This allows us to describe the reasons behind the model’s decisions and highlight the distinctions between classes.
@@
Feature Type Feature Type calculation
name_len count screen_name_sim real-valued Jaccard of (screen_name, user_name)
screen_name_len count foll_friends real-valued follower / friends
description_len count age real-valued created_at / collection date
listed count listed_by_age real-valued listed / account age
statuses count statuses_by_age real-valued statuses / account age
followers count followers_by_age real-valued followers / account age
following count following_by_age real-valued friends / account age
name_upper_len count name_upper_pcnt real-valued percentage of upper case
name_lower_len count name_lower_pcnt real-valued percentage of lower case
name_digits_len count name_digits_pcnt real-valued percentage of digits
name_special_len count name_special_pcnt real-valued percentage of other characters
screen_name_upper_len count screen_name_upper_pcnt real-valued percentage of upper case
screen_name_lower_len count screen_name_lower_pcnt real-valued percentage of lower case
screen_name_digits_len count screen_name_digits_pcnt real-valued percentage of digits
screen_name_special_len count screen_name_special_pcnt real-valued percentage of other characters
description_upper_len count description_upper_pcnt real-valued percentage of upper case
description_lower_len count description_lower_pcnt real-valued percentage of lower case
description_digits_len count description_digits_pcnt real-valued percentage of digits
description_special_len count description_special_pcnt real-valued percentage of other characters
description_urls count name_entropy real-valued entropy of user name
description_mentions count screen_name_entropy real-valued entropy of the screen name
description_hashtags count has_location boolean
total_urls count has_profile_image boolean
protected boolean has_profile_url boolean
verified boolean
@@
Method Type C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TB-20 TB-22 Average
SGBot F 77.9 72.1 94.6 99.5 82.3 82.7 49.6 84.9 36.6 75.57
Kudugunta et al. F 75.3 49.8 91.7 94.5 50.9 49.2 49.6 47.3 51.7 62.22
Hayawi et al. F 85.6 34.7 93.8 91.5 60.8 60.9 20.5 77.1 24.7 61.06
BotHunter F 97.2 69.2 91.6 99.6 82.2 82.9 49.6 79.1 23.5 74.98
NameBot F 83.4 44.8 85.7 91.6 61.1 67.5 38.5 65.1 0.5 59.80
Abreu et al. F 76.4 66.7 95.0 97.9 76.9 83.5 53.8 77.1 53.4 75.63
BotArtist F 98.3 76.1 97.0 99.7 80.6 88.3 68.4 82.2 58.2 83.19
Cresci T 1.17 - 22.8 - - - - 13.7 - -
Wei T 82.7 - 78.4 - - - - 57.3 53.6 -
BGSRD T 90.8 35.7 86.3 90.5 58.2 41.1 13.0 70.0 21.1 56.30
RoBERTa T 95.8 - 94.3 - - - - 73.1 20.5 -
T5 T 89.3 - 92.3 - - - - 70.5 20.2 -
Efthimion FT 94.1 5.2 91.8 95.9 68.2 71.7 0.0 67.2 27.5 57.95
Kantepe FT 78.2 - 79.4 - - - - 62.2 58.7 -
Miller FT 83.8 59.9 86.8 91.1 56.8 43.6 0.0 74.8 45.3 60.23
Varol FT 94.7 - - - - - - 81.1 27.5 -
Kouvela FT 98.2 66.6 99.1 98.2 80.4 81.1 28.1 86.5 30.0 74.24
Santos FT 78.8 14.5 83.0 92.4 65.2 75.7 21.0 60.3 - -
Lee FT 98.6 67.8 99.3 97.9 82.5 82.7 50.3 80.0 30.4 76.61
LOBO FT 98.8 - 97.7 - - - - 80.8 38.6 -
Moghaddam FG 73.9 - - - - - - 79.9 32.1 -
Alhosseini FG 92.2 - - - - - - 72.0 38.1 -
Knauth FTG 91.2 39.1 93.4 91.3 94.0 54.2 41.3 85.2 37.1 69.64
FriendBot FTG 97.6 - 87.4 - - - - 80.0 - -
SATAR FTG 95.0 - - - - - - 86.1 - -
Botometer FTG 66.9 77.4 96.1 46.0 79.6 79.0 30.8 53.1 42.8 63.5
Rodrifuez-Ruiz FTG 87.7 - 85.7 - - - - 63.1 56.6 -
GraphHist FTG 84.5 - - - - - - 67.6 - -
EvolveBot FTG 90.1 - - - - - - 69.7 14.1 -
Dehghan FTG 88.3 - - - - - - 76.2 - -
GCN FTG 97.2 - - - - - - 80.8 54.9 -
GAT FTG 97.6 - - - - - - 85.2 55.8 -
HGT FTG 96.9 - - - - - - 88.2 39.6 -
SimpleHGN FTG 97.5 - - - - - - 88.2 45.4 -
BotRGCN FTG 97.3 - - - - - - 87.3 57.5 -
RGT FTG 97.8 - - - - - - 88.0 42.9 -
@@
Table 3: The performance of each selected bot detection model, as reported in the (Feng et al. 2021b) paper, is compared with that of BotArtist. Performance is measured using the F1-score. In this benchmark, each model is trained and tested on each dataset separately.
Method C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TB-20 TB-22 Total Average
BotArtist 82.7 39.9 87.3 99.0 80.6 73.8 16.6 80.3 56.9 63.7 68.5
Lee 82.3 0.0 83.6 97.7 78.2 67.7 20.0 8.5 42.4 52.9 53.3
Abreu 84.4 0.3 80.1 88.4 67.1 40.8 11.7 15.6 29.0 40.4 46.3
SGBot 75.0 3.6 79.8 99.2 76.7 68.9 0.0 15.2 43.3 53.8 51.5
BotHunter 73.4 7.1 76.0 99.2 76.0 44.8 11.1 14.7 28.0 43.1 47.8
Kouvela 95.5 20.4 94.7 98.1 78.4 71.4 21.0 28.5 36.0 52.0 60.5
Botometer 66.9 77.4 96.1 46.0 79.6 79.0 30.8 53.1 42.8 45.3 63.5
@@
Table 4: The measurement of usage, like general bot detection approaches, involves training and testing models on all datasets. Performance is assessed using the F1-score.
Feature Extraction
@@
As mentioned earlier, in this study we pursue the creation of a simple model that in comparison with other bot detection systems does not require a large volume of information or features for accurate prediction. In our case, the number of Twitter profiles triggers our attention. The majority of social media bot accounts are created for content promotion and increasing the attention of other registered users. To maximize this goal, bot accounts pursue the expansion of social audience. Social expansion is very difficult to achieve without high activity, such as tweets, re-tweets, comments, etc. In comparison with profile-based features, the extraction and utilization of textual or graph features can result in instability, causing a significant performance drop across various periods or discussion topic datasets (Shevtsov et al. 2023a). Consequently, we need to extract as much information as possible from the user profiles to identify extremely active accounts. The difference between very popular social media accounts (also known as celebrities), and bot accounts is that they try to post and share as much content as possible during the day. To properly separate normal and bot accounts, we develop 49 unique features described in table: 2. Due to the differences between Twitter API v1.1 and v2, we have devised a method for constructing features in a way that allows the extraction of selected information from both v1.1 and v2 without any additional manipulation. The profile features we have extracted are categorized into three distinct groups: count, boolean, and real-valued.
@@
The initial count feature categories encompass raw Twitter API object values, including followers, followings, status, and number of subscribed Twitter lists. In addition to these raw values, we have calculated statistics for textual fields such as user name, screen_name, and description. For these features, we have quantified the number of distinct characters, including uppercase letters, lowercase letters, digits, and special characters that do not fall into any of the previous categories. Furthermore, we have leveraged the profile description field, where users can provide additional information such as mentions of other users/organizations, hashtags, and additional URLs. In this context, we have counted the volume of the mentioned entities in the description field as raw values.
@@
In contrast to the count categories, where values are represented as integers, the real-valued category contains floating-point values. We have conducted various comparisons and measurements across the profile fields in this category. Initially, we compute the age of the accounts, which allows us to distinguish highly active users with millions of posts spanning several years from users with similar activity but with an active period of only a few days. Account age is measured in days, calculated as the duration between the account creation date, provided by the Twitter API, and the collection date. Based on the account’s age, we can compute metrics similar to those in the count values category but adjusted for the account’s age perspective. For example, we compute metrics such as statuses_by_age, follower_by_age, and following_by_age using the formula:
$feature\_by\_age=\frac{feature\ value}{profile\ age}$
@@
Given that many bot accounts are mass-created by automated scripts, they often have very similar or even trivial user names and screen names. Therefore, we have also applied the Jaccard similarity measure between user names and screen names. Jaccard similarity calculates the intersection of characters used in both fields in relation to the differences between them, using the following formula:
$Jaccard\ similarity(A,B)=\frac{|A\cap B|}{|A\cup B|}$
@@
Additionally, we have computed the entropy of the user name and screen_name, enabling us to assess the randomness of these selected fields. In addition to the count values derived from the profile text fields (user name, screen_name, and description), we have computed the percentage of each specific character category, including lowercase, uppercase, digits, and special characters, relative to the length of the text field. These measurements offer a more precise means of identifying textual preferences more flexible to varying text lengths since they are computed as percentages of the field length.
@@
By employing this comprehensive feature engineering approach, we have maximized feature extraction while working with the limited profile information available through Twitter API v2. In total, we have extracted 49 unique profile features, representing the full set of features used before the feature selection process in the BotArtist model.
Experimental Results
@@
According to the methodology of SAMLP described earlier, we’ve successfully developed a semi-automated machine learning pipeline capable of constructing classification models (both binary and multi-class) for both balanced and imbalanced datasets. Using this pipeline, we created and tested our custom implementation called BotArtist for Twitter bot detection, utilizing nine well-known datasets. To evaluate BotArtist’s performance, in comparison to 34 other existing research approaches, we employed the most comprehensive Twitter bot detection benchmark available (Feng et al. 2022b), where various methods are compared using identical training/validation and testing data portions. This bench-marking allowed us to make a fair comparison between our approach and established bot detection methods.
@@
In our comparison, we not only evaluate BotArtist against bot detection approaches based on similar categories of features but also compare it to existing approaches developed based on different sets of characteristics. The selected methods are categorized into five different groups.
Profile-Based features (F)
This category includes approaches that rely on features extracted from user profile objects, including SGBot (Yang et al. 2020), Kudugunta (Kudugunta and Ferrara 2018), Hayawi (Hayawi et al. 2022), BotHunter (Beskow and Carley 2018), NameBot (Beskow and Carley 2019) and Abreu (Abreu, Ralha, and Gondim 2020) implementations.
Text-Based features (T)
In this category, we have approaches that focus on text characteristics, including Cresci (Cresci et al. 2016), Wei (Wei and Nguyen 2019), BGSRD (Guo et al. 2021a), RoBERTa (Liu et al. 2019) and T5 (Raffel et al. 2020).
Profile and Text-Based features (FT)
These approaches combine both profile and text-based features, and the selected implementations are Efthimion (Efthimion, Payne, and Proferes 2018), Kantepe (Kantepe and Ganiz 2017), Miller (Miller et al. 2014), Varol (Varol et al. 2017), Kouvela (Kouvela, Dimitriadis, and Vakali 2020), Lee (Lee, Eoff, and Caverlee 2011) and LOBO (Echeverria et al. 2018).
Profile and Graph-Based features (FG)
Some approaches combine profile features with graph-based information. This category includes Moghaddam (Moghaddam and Abbaspour 2022) and Alhosseini (Ali Alhosseini et al. 2019) implementations.
Profile, Text, and Graph-Based features (FTG)
Approaches in this category aim to extract and combine as many features as possible, encompassing profile, text, and graph-based information. The selected implementations are Knauth (Knauth 2019), FriendBot (Beskow and Carley 2020), SATAR (Feng et al. 2021a), Botometer (Yang, Ferrara, and Menczer 2022), Rodríguez-Ruiz (Rodríguez-Ruiz et al. 2020), GraphHist (Magelinski, Beskow, and Carley 2020), EvolveBot (Yang, Harkreader, and Gu 2013), Dehghan (Dehghan et al. 2023), GCN (Kipf and Welling 2016), GAT (Veličković et al. 2017), HGT (Hu et al. 2020), SimpleHGN (Lv et al. 2021), BotRGCN (Feng et al. 2021c) and RGT (Feng et al. 2022a).
Models comparison
@@
Our goal with BotArtist was to create a model with the highest possible level of generalizability. To achieve this, we designed two separate comparison scenarios. Initially, our focus was on assessing the fine-tuned BotArtist’s ability to capture general patterns within each dataset in our collection. As a result, we conducted training and testing for each of the selected models, including BotArtist, on each of the nine chosen datasets individually. This comparative approach allowed us to evaluate how well these selected models performed on each specific dataset, serving as a more rigorous bot detection implementation.
@@
Additionally, the designed experiments provided us with a broader understanding of the overall capabilities of the selected methods when applied in more general scenarios. While implementing these experiments, we took into account that the performance would likely be higher than that of a generic approach. Therefore, our primary focus was on discerning the differences between the models to select the most accurate bot detection methods for the broader, more generalized comparison.
@@
Previously, we had evaluated and compared models on each isolated dataset and calculated their average performance. While this approach provided us with an initial insight into model performance across diverse datasets, it lacked an assessment of model generalizability. To address this, we designed additional comparisons where we assessed the model’s ability to capture general patterns across different datasets. In these cases, we combined the training data from the nine datasets into a single dataset for model training. Combining data from multiple datasets, each with varying periods, discussion topics, and communities, demanded a high level of model generalization to avoid overfitting.
@@
To measure the performance of the trained methods, we evaluated them by testing on portions of each dataset and also on the union of all the dataset testing portions. These approaches allowed us to not only identify the general performance of each bot detection method but also gain insights into performance variations across each specific dataset.
Parameter Value
max_depth
learning_rate
subsample
colsample_bytree
min_child_weight
gamma
reg_lambda
n_estimators
eval_metric
tree_method
predictor
objective
scale_pos_weight
selected features
Table 5: The following are the selected parameters of the final general BotArtist (XGBoost) model, along with the number of features chosen during the feature selection procedure.
In the initial model comparison experiment, we compared selected methods and assessed their performance based on the F1-score. This choice of metric was motivated by the highly imbalanced nature of the datasets, where proper comparison necessitates measurement metrics that account for class imbalances.
In table: 3, we present the results of this comparison. Notably, the BotArtist model emerged as the only bot detection method that achieved superior performance across three different datasets (M-18, C-R-19, B-F-19). Furthermore, the BotArtist model demonstrated the highest average F1-score, reaching 83.19. This represents a substantial improvement of 6.5% over the most accurate of the existing methods, showcasing the model’s capacity to accurately capture patterns within each distinct dataset.
As shown in table: 3, some datasets posed challenges for more complex methods that rely on text and graph features. These methods struggled to utilize datasets without text or graph information due to their limitations in accessing essential information required for the model.
In the second model comparison scenario, our objective was to assess the generalization capabilities of various models. To achieve this, we selectively choose the most accurate models from table: 3, in addition to the Botometer model (Yang, Ferrara, and Menczer 2022). The Botometer model was included because it had already been trained on a diverse range of datasets and was expected to perform well in general-case scenarios. Additionally, the SAMLP pipeline manages to fine-tune BotArtists with the parameters presented in the table: 5 where the XGBoost classification model achieves higher performance during the K-Fold cross-validation.

### v4

@@ 4.2 Model Fine-Tuning
Having reduced the problem’s dimensionality through feature selection, we now focus on identifying the model and its configuration (hyper-parameter set) that can provide accurate predictions over unseen data samples. We select three well-known machine learning classification models: SVM (Boser et al., 1992), RandomForest (Breiman, 2001), and state-of-the-art XGBoost (Extreme Gradient Boosting) (Chen et al., 2015).
@@
For each model, we define large hyper-parameter ranges to cover a variety of configurations. To reduce complexity and execution time, we develop a sampling method that randomly selects $C$ unique configurations per model (in our case, $C=50$). This method allows us to estimate different possible configurations without sacrificing execution time. Additionally, since the dataset may be affected by class imbalance, we compute class-specific weights to construct selected classifiers with consideration of class imbalance.
@@
Afterward, we select the best classifier and configuration based on the average validation (F1-score). This model configuration becomes the final model, which will be evaluated over the hold-out (testing) data portion. In the final stage of the model development, we provide model explainability using the SHAP game-theoretical approach (Shapley, 1953). This allows us to describe the reasons behind the model’s decisions and highlight the distinctions between classes.
@@
Feature Type Feature Type Calculation name_len count screen_name_sim real-valued Jaccard of (screen_name, user_name) screen_name_len count foll_friends real-valued follower / friends description_len count age real-valued created_at / collection date listed count listed_by_age real-valued listed / account age statuses count statuses_by_age real-valued statuses / account age followers count followers_by_age real-valued followers / account age following count following_by_age real-valued friends / account age name_upper_len count name_upper_pcnt real-valued percentage of upper case name_lower_len count name_lower_pcnt real-valued percentage of lower case name_digits_len count name_digits_pcnt real-valued percentage of digits name_special_len count name_special_pcnt real-valued percentage of other characters screen_name_upper_len count screen_name_upper_pcnt real-valued percentage of upper case screen_name_lower_len count screen_name_lower_pcnt real-valued percentage of lower case screen_name_digits_len count screen_name_digits_pcnt real-valued percentage of digits screen_name_special_len count screen_name_special_pcnt real-valued percentage of other characters description_upper_len count description_upper_pcnt real-valued percentage of upper case description_lower_len count description_lower_pcnt real-valued percentage of lower case description_digits_len count description_digits_pcnt real-valued percentage of digits description_special_len count description_special_pcnt real-valued percentage of other characters description_urls count name_entropy real-valued entropy of user name description_mentions count screen_name_entropy real-valued entropy of the screen name description_hashtags count has_location boolean total_urls count has_profile_image boolean protected boolean has_profile_url boolean verified boolean
@@
Table 3: The performance of each selected bot detection model, as reported in the (Feng et al., 2021c) paper, is compared with that of BotArtist. Performance is measured using the F1-score. In this benchmark, each model is trained and tested on each dataset separately.
Method Type C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TB-20 TB-22 Average SGBot F 77.9 72.1 94.6 99.5 82.3 82.7 49.6 84.9 36.6 75.57 Kudugunta et al. F 75.3 49.8 91.7 94.5 50.9 49.2 49.6 47.3 51.7 62.22 Hayawi et al. F 85.6 34.7 93.8 91.5 60.8 60.9 20.5 77.1 24.7 61.06 BotHunter F 97.2 69.2 91.6 99.6 82.2 82.9 49.6 79.1 23.5 74.98 NameBot F 83.4 44.8 85.7 91.6 61.1 67.5 38.5 65.1 0.5 59.80 Abreu et al. F 76.4 66.7 95.0 97.9 76.9 83.5 53.8 77.1 53.4 75.63 BotArtist F 98.3 76.1 97.0 99.7 80.6 88.3 68.4 82.2 58.2 83.19 Cresci T 1.17 - 22.8 - - - - 13.7 - - Wei T 82.7 - 78.4 - - - - 57.3 53.6 - BGSRD T 90.8 35.7 86.3 90.5 58.2 41.1 13.0 70.0 21.1 56.30 RoBERTa T 95.8 - 94.3 - - - - 73.1 20.5 - T5 T 89.3 - 92.3 - - - - 70.5 20.2 - Efthimion FT 94.1 5.2 91.8 95.9 68.2 71.7 0.0 67.2 27.5 57.95 Kantepe FT 78.2 - 79.4 - - - - 62.2 58.7 - Miller FT 83.8 59.9 86.8 91.1 56.8 43.6 0.0 74.8 45.3 60.23 Varol FT 94.7 - - - - - - 81.1 27.5 - Kouvela FT 98.2 66.6 99.1 98.2 80.4 81.1 28.1 86.5 30.0 74.24 Santos FT 78.8 14.5 83.0 92.4 65.2 75.7 21.0 60.3 - - Lee FT 98.6 67.8 99.3 97.9 82.5 82.7 50.3 80.0 30.4 76.61 LOBO FT 98.8 - 97.7 - - - - 80.8 38.6 - Moghaddam FG 73.9 - - - - - - 79.9 32.1 - Alhosseini FG 92.2 - - - - - - 72.0 38.1 - Knauth FTG 91.2 39.1 93.4 91.3 94.0 54.2 41.3 85.2 37.1 69.64 FriendBot FTG 97.6 - 87.4 - - - - 80.0 - - SATAR FTG 95.0 - - - - - - 86.1 - - Botometer FTG 66.9 77.4 96.1 46.0 79.6 79.0 30.8 53.1 42.8 63.5 Rodrifuez-Ruiz FTG 87.7 - 85.7 - - - - 63.1 56.6 - GraphHist FTG 84.5 - - - - - - 67.6 - - EvolveBot FTG 90.1 - - - - - - 69.7 14.1 - Dehghan FTG 88.3 - - - - - - 76.2 - - GCN FTG 97.2 - - - - - - 80.8 54.9 - GAT FTG 97.6 - - - - - - 85.2 55.8 - HGT FTG 96.9 - - - - - - 88.2 39.6 - SimpleHGN FTG 97.5 - - - - - - 88.2 45.4 - BotRGCN FTG 97.3 - - - - - - 87.3 57.5 - RGT FTG 97.8 - - - - - - 88.0 42.9 -
@@
Table 4: The measurement of performance in the case of general bot detection approaches, involves training and testing models on all datasets. Performance is assessed using the F1-score.
Method C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TB-20 TB-22 Total Average BotArtist 82.7 39.9 87.3 99.0 80.6 73.8 16.6 80.3 56.9 63.7 68.5 Lee 82.3 0.0 83.6 97.7 78.2 67.7 20.0 8.5 42.4 52.9 53.3 Abreu 84.4 0.3 80.1 88.4 67.1 40.8 11.7 15.6 29.0 40.4 46.3 SGBot 75.0 3.6 79.8 99.2 76.7 68.9 0.0 15.2 43.3 53.8 51.5 BotHunter 73.4 7.1 76.0 99.2 76.0 44.8 11.1 14.7 28.0 43.1 47.8 Kouvela 95.5 20.4 94.7 98.1 78.4 71.4 21.0 28.5 36.0 52.0 60.5 Botometer 66.9 77.4 96.1 46.0 79.6 79.0 30.8 53.1 42.8 45.3 63.5
4.3 Feature Extraction
@@
As mentioned earlier, in this study we pursue the creation of a simple model that in comparison with other bot detection systems- does not require a large volume of information or features for accurate prediction. In our case, the number of Twitter profiles triggers our attention. The majority of social media bot accounts are created for content promotion and increasing the attention of other registered users. To maximize this goal, bot accounts aim for the expansion of social audience. Social expansion is very difficult to achieve without high activity, such as tweets, re-tweets, comments, etc. In comparison with profile-based features, the extraction and utilization of textual or graph features can result in instability, causing a significant performance drop across various periods or discussion topic datasets (Shevtsov et al., 2024). Consequently, we need to extract as much information as possible from the user profiles to identify extremely active accounts. The difference between very popular social media accounts (also known as celebrities), and bot accounts is that they try to post and share as much content as possible during the day. Due to the differences between Twitter API v1.1 and v2, we have devised a method for constructing features in a way that allows the extraction of selected information from both v1.1 and v2 without any additional manipulation. The profile features we have extracted are categorized into three distinct groups: count, boolean, and real-valued.
@@
The initial count feature categories encompass raw Twitter API object values, including followers, followings, status, and number of subscribed Twitter lists. In addition to these raw values, we calculate statistics for textual fields such as user name, screen_name, and description. For these features, we quantify the number of distinct characters, including uppercase letters, lowercase letters, digits, and special characters that do not fall into any of the previous categories. Furthermore, we leverage the profile description field, where users can provide additional information such as mentions of other users/organizations, hashtags, and additional URLs. In this context, we have counted the volume of the mentioned entities in the description field as raw values.
@@
In contrast to the count categories, where values are represented as integers, the real-valued categories contain floating-point values. We conduct various comparisons and measurements across the profile fields in this category. Initially, we compute the age of the accounts, which allows us to distinguish highly active users with millions of posts spanning several years from users with similar activity but with an active period of only a few days. Account age is measured in days, calculated as the duration between the account creation date, provided by the Twitter API, and the collection date. Based on the account’s age, we can compute metrics similar to those in the count values category but adjusted for the account’s age perspective.
@@
Given that many bot accounts are mass-created by automated scripts, they often have very similar or even trivial user names and screen names. Therefore, we apply the Jaccard similarity measure between user names and screen names to calculate the intersection of characters used in both fields to the differences between them.
@@
Additionally, we compute the entropy of the user name and screen_name, enabling us to assess the randomness of these selected fields. In addition to the count values derived from the profile text fields (user name, screen_name, and description), we compute the percentage of each specific character category, including lowercase, uppercase, digits, and special characters, relative to the length of the text field. These measurements offer a more precise means of identifying textual preferences more flexible to varying text lengths since they are computed as percentages of the field length.
@@
By employing this comprehensive feature engineering approach, we maximize the feature extraction while working with the limited profile information available through Twitter API v2. In total, we extract 49 unique profile features Table 2, representing the full set of features used before the feature selection process in the BotArtist model.
5 Experimental Results
@@
According to the methodology of SAMLP described earlier, we successfully develop a semi-automated machine learning pipeline capable of constructing classification models (both binary and multiclass) for both balanced and imbalanced datasets. Using this pipeline, we create and test our custom implementation called BotArtist for Twitter bot detection, utilizing nine well-known datasets. To evaluate BotArtist’s performance, in comparison to 35 other existing research approaches, we employ the most comprehensive Twitter bot detection benchmark available (Feng et al., 2022a), where various methods are compared using identical training/validation and testing data portions. This benchmarking allows us to make a fair comparison between our approach and established bot detection methods.
@@
In our comparison, we not only evaluate BotArtist against bot detection approaches based on similar categories of features but also compare it to existing approaches developed based on different sets of characteristics. The selected methods are categorized into five different groups: feature-based (F), textual (T), graph (G), and any combination of them.
5.1 Models comparison
@@
Our goal with BotArtist is to create a model with the highest possible level of generalizability. To achieve this, we design two separate comparison scenarios. Initially, our focus is on assessing the fine-tuned BotArtist’s ability to capture general patterns within each dataset in our collection. As a result, we conduct training and testing for each of the selected models, including BotArtist, on each of the nine chosen datasets individually. This comparative approach allows us to evaluate how well these selected models perform on each specific dataset, serving as a more rigorous bot detection implementation.
@@
Additionally, the designed experiments provide us with a broader understanding of the overall capabilities of the selected methods when applied in more general scenarios. While implementing these experiments, we take into account that the performance would likely be higher than that of a generic approach. Therefore, our primary focus is on discerning the differences between the models to select the most accurate bot detection methods for the broader, more generalized comparison. Based on the designed experiment we identified that (see Table 3) the BotArtist model emerged as the only bot detection method that achieves superior performance across three different datasets (M-18, C-R-19, B-F-19). Furthermore, the BotArtist model demonstrates the highest average F1-score, reaching 83.19. This represents a substantial improvement of 6.5% over the most accurate of the existing methods. As shown in Table 3, some datasets posed challenges for more complex methods that rely on text and graph features. These methods struggle to utilize datasets without text or graph information due to their limitations in accessing essential information required for the model.
@@
Previously, we had evaluated and compared models on each isolated dataset and calculated their average performance. While this approach provided us with an initial insight into model performance across diverse datasets, it lacked an assessment of model generalizability. To address this, we design additional comparisons where we assess the model’s ability to capture general patterns across different datasets. In these cases, we combine the training data from the nine datasets into a single dataset for model training. Combining data from multiple datasets, each with varying periods, discussion topics, and communities, demands a high level of model generalization to avoid overfitting.
@@
To measure the performance of the trained methods, we evaluate them by testing on portions of each dataset and also on the union of all the dataset testing portions. These approaches allow us to not only identify the general performance of each bot detection method but also gain insights into performance variations across each specific dataset. To achieve this, we selectively choose the most accurate models from Table 3, in addition to the Botometer model (Yang et al., 2022). The Botometer model is included because it has already been trained on a diverse range of datasets and is expected to perform well in general-case scenarios.